# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ 4.3 Simple examples
For the examples from subsection 2.3, both the basic version and its variation, we provide a reference choice structure, and a list of adapted choices, one for each exogenous information structure from example 3.3. Recall that $\Omega=\{1,2\}$ and $\mathscr{A}=\mathcal{P}(\Omega)$.
@@
$\mathscr{C}=(\mathscr{C}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ defines a reference choice structure on $(F,\pi,{\bm{\mathsf{X}}})$.
@@ 1. $c_{k\bullet}$ : $k\in\{1,2\}$ $c_{km},c_{\bullet m}$ : $k,m\in\{1,2\}$
2.(b) $c_{k\bullet}$ : $k\in\{1,2\}$ $c_{1g},c_{2m},c_{\bullet m}$ : $k,m\in\{1,2\},g\in M$
2.(c) $c_{k\bullet}$ : $k\in\{1,2\}$ $c_{1m},c_{2g},c_{\bullet m}$ : $k,m\in\{1,2\},g\in M$
@@ 3. $c_{f\bullet}$ : $f\in M$ $c_{kg},c_{\bullet g}$ : $k\in\{1,2\},g\in M$
$c^{\prime}_{kg}=$$\leavevmode\nobreak\ \{(\omega,k^{\prime},m^{\prime})\in W\mid k^{\prime}=k,\leavevmode\nobreak\ m^{\prime}=g(\omega)\},$
@@
$\mathscr{C}^{\prime}=(\mathscr{C}^{\prime}_{{\bm{\mathsf{x}}}^{\prime}})_{{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime}}$ defines a reference choice structure on $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$.
@@ 3. $c^{\prime}_{f\bullet}$ : $f\in M$ $c^{\prime}_{kg},c^{\prime}_{\bullet g}$ : $k\in\{1,2\},g\in M$
Also note that a choice like $c_{kg}$ is conditional on the knowledge that ${\bm{\mathsf{x}}}_{k}$ has been chosen at the root of ${\bm{\mathsf{X}}}$, while $c_{\bullet g}$ is independent of the initial choice. This is reflected by the respective sets of immediate predecessors, see lemma A.44. A similar remark is true for the variant, see lemma A.45. Hence, as in [3, chapter 5], choices can reflect the endogenous information, that is, the information about the position in the decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$, agents have. While $c_{kg}$ is a choice of perfect (endogenous) information, $c_{\bullet g}$ is not. The discussion on this theme will be continued in the second paper.
@@ 4.4 Action path stochastic decision forests
Finally, we consider the action-path sdf $(F,\pi,{\bm{\mathsf{X}}})$ from subsection 2.4 over some exogenous scenario space $(\Omega,\mathscr{A})$. Let $W=\bigcup F$.
Let $t\in\mathbb{T}$. For any set $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and any family $A_{t}=(A_{t,\omega})_{\omega\in\Omega}$ of Borel sets in $\mathbb{A}$, let
$c(A_{<t},A_{t})=\{(\omega,f)\in W\mid f|_{[0,t)_{\mathbb{T}}}\in A_{<t},\leavevmode\nobreak\ f(t)\in A_{t,\omega}\}.$
@@
So we consider choices that correspond to actions at a predefined time $t$. Again, there is some sort of duality here: action paths are the result of progressive choosing; choices are collections of action paths, essentially. More precisely, $c(A_{<t},A_{t})\in\mathscr{C}_{t}$ describes the choice of an action in $A_{t,\omega}$ given scenario $\omega$ and at time $t$, given the history is contained in $A_{<t}$. We assume that such an action is really possible for at least some scenario (AP.C0), that it really constitutes a choices in that there is an alternative (AP.C1), and that it is complete in the sense that it only trivially intersects with random moves (AP.C2).
@@
$\downarrow c=\{x_{u}(w)\mid w\in c,\leavevmode\nobreak\ u\in\mathbb{T}\colon t<u\}\cup\{\{w\}\mid w\in c\}.$
@@
For the remainder of this subsection, we impose assumption AP.W4, see subsection 2.4. We fix both an agent $i\in I$ and an exogenous information structure $\mathscr{F}^{i}=(\mathscr{F}^{i}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{x}}}}$ this agent is equipped with. We denote the canonical projection $\mathbb{A}\to\mathbb{A}^{i}$ by $p^{i}$. For all $t\in\mathbb{T}$ and ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, let $\mathscr{C}^{i}_{{\bm{\mathsf{x}}}}$ the set of all sets $c(A_{<t},A_{t})$ as above such that
@@
    $A_{t}=(A_{t,\omega})_{\omega\in\Omega}$ such that there is $A_{t}^{i}\in\mathscr{B}(\mathbb{A}^{j})$ satisfying, for all $\omega\in\Omega$,
@@
Hence, $c(A_{<t},A_{t})$ allows for choosing a measurable set of individual actions for agent $i$ at the random move ${\bm{\mathsf{x}}}$.
@@
$\mathscr{C}^{i}=(\mathscr{C}^{i}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ defines a reference choice structure.
Let $t\in\mathbb{T}$, $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$, $D\in\mathscr{A}$, and $g\colon D\to\mathbb{A}^{i}$. Let $A_{t}^{i,g}=(A_{t,\omega}^{i,g})_{\omega\in\Omega}$ be given by
@@
Let $c(A_{<t},i,g)=c(A_{<t},A_{t}^{i,g})$. Provided it is an element of $\mathscr{C}_{t}$, this models the choice of agent $i$, given a history in $A_{<t}$, to take the action $g(\omega)$ in scenario $\omega\in D$ at time $t$. The perfect (endogenous) information case corresponds to $A_{<t}=\{f|_{[0,t)_{\mathbb{T}}}\}$ for some $f\in\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f}\neq\emptyset$.
@@
The next and final theorem of this paper is concerned with the following question: Provided $c=c(A_{<t},i,g)\in\mathscr{C}_{t}$, what is the link between the adaptedness of $c$ and the measurability of the function $g|_{D_{\bm{\mathsf{x}}}}$ with respect to $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$, for all ${\bm{\mathsf{x}}}$ that $c$ is available at? [^14] As stated in the theorem below, the availability of $c$ at ${\bm{\mathsf{x}}}$ implies $D_{\bm{\mathsf{x}}}\subseteq D$. Indeed, in stochastic control and game theory, action path specifications are often implicitly used and choices are defined by assigning an action $g(\omega)$ to any scenario $\omega$, for $g$ measurable with respect to the value of the underlying filtration at the current time (see, e.g., [10, chapter 8], [34, 14, 13]). Hence, if we can answer the question above by providing a tight link, then this customary modelling paradigm in stochastic control and game theory can be explained as a derivative of adapted choices in stochastic decision forests and can therefore be interpreted in terms of traditional decision theory.
@@
The theorem below affirms that the $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurability of $g|_{D_{\bm{\mathsf{x}}}}$ for relevant ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ is sufficient for the adaptedness of $c$. Moreover, it is also necessary, provided the reference choice structure sufficiently reflects the Borel $\sigma$-algebra on $\mathbb{A}^{i}$. Precisely, we consider the following assumption:
  • 1.
    Assumption AP.C3. For all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ such that, with $t=\mathfrak{t}({\bm{\mathsf{x}}})$, there are $D\in\mathscr{A}$, $g\colon D\to\mathbb{A}^{i}$ and $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ satisfying $\tilde{c}=c(A_{<t},i,g)\in\mathscr{C}_{t}$ and making $\tilde{c}$ available at ${\bm{\mathsf{x}}}$, there are
    1. (a)
      $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ such that for all $(\omega,f)\in\tilde{c}$, $f|_{[0,t)_{\mathbb{T}}}\in A^{\prime}_{<t}$, and
    2. (b)
@@
      an intersection-stable generator $\mathscr{G}(\mathbb{A}^{i})$ of the Borel $\sigma$-algebra of $\mathbb{A}^{i}$ such that for all $G\in\mathscr{G}(\mathbb{A}^{i})$, upon letting $A^{i,G}_{t}=(A^{i,G}_{t,\omega})_{\omega\in\Omega}$ be given by $A^{i,G}_{t,\omega}=(p^{i})^{-1}(G)$ for $\omega\in D_{\bm{\mathsf{x}}}$ and $A^{i,G}_{t,\omega}=\emptyset$ for $\omega\notin D_{\bm{\mathsf{x}}}$, we have
      $c(A^{\prime}_{<t},A^{i,G}_{t})\in\mathscr{C}^{i}_{{\bm{\mathsf{x}}}}\cup\{\emptyset\}.$
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be the action path sdf over $(\Omega,\mathscr{A})$ such that $W=\bigcup F$ satisfies assumption AP.W4. Further, let $i\in I$, $\mathscr{F}^{i}$ be an exogenous information structure and $\mathscr{C}^{i}$ be the reference choice structure defined above.
Furthermore, let $t\in\mathbb{T}$, $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$, $D\in\mathscr{A}$, and $g\colon D\to\mathbb{A}^{i}$ such that $c(A_{<t},i,g)\in\mathscr{C}_{t}$.
Then $c(A_{<t},i,g)$ is a non-redundant and complete choice. Moreover, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ that $c(A_{<t},i,g)$ is available at, we have $D_{\bm{\mathsf{x}}}\subseteq D$, and the following implications hold true:
@@
    If $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}}$-measurable, then $c(A_{<t},i,g)$ is $\mathscr{F}^{i}$-$\mathscr{C}^{i}$-adapted.
@@
    If assumption AP.C3 is satisfied and $c(A_{<t},i,g)$ is $\mathscr{F}^{i}$-$\mathscr{C}^{i}$-adapted, then $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}}$-measurable.
@@
  • 1.
    For $\mathbb{T}=\mathbb{R}_{+}$, singleton $\Omega$, singleton $A_{<t}=\{f|_{[0,t)_{\mathbb{T}}}\}$ for $t\in\mathbb{T}$, and $W=\Omega\times\mathbb{A}^{\mathbb{T}}\cong\mathbb{A}^{\mathbb{T}}$, $c(A_{<t},i,g)$ corresponds to the so-called “differential game” choice in [2, section 2] (see also [5, example 4.14]), where agent $i$ chooses the value of $g$ given the history $f|_{[0,t)_{\mathbb{T}}}$. In that sense, the adapted choices constructed in theorem 4.40 generalise this example in several respects, most importantly along the stochastic dimension. If $\mathbb{T}=\{0,1,\dots,N\}$ for some $N\in{\mathbb{N}^{\ast}}$ or $\mathbb{T}=\mathbb{N}$, then we get a model of the usual specification of choices from decision problems in discrete time (see, e.g., [10]).
  • 2.
@@
  • 3.
    If $W=\Omega\times\mathbb{A}^{\mathbb{T}}$, and $\mathbb{A}^{i}$ has at least two elements, and $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ is non-empty for some $t\in\mathbb{T}$, then for any map $g\colon\Omega\to\mathbb{A}^{i}$ we have $c(A_{<t},i,g)\in\mathscr{C}_{t}$, that is, it satisfies assumptions AP.C0, AP.C1, and AP.C2. Moreover, assumption AP.C3 is satisfied.
  • 4.
    In case of the timing game (see example 2.23, and, e.g., [40, 29]), we have $\mathbb{A}^{i}=\{0,1\}$. Let $t\in\mathbb{T}$. If $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ is a non-empty set of component-wise decreasing paths $f_{t}$ such that $p^{i}\circ f_{t}=1_{[0,t)_{\mathbb{T}}}$, then, for all $g\colon\Omega\to\{0,1\}$, we have $c(A_{<t},i,g)\in\mathscr{C}_{t}$, that is, it satisfies assumptions AP.C0, AP.C1, AP.C2. Moreover, assumption AP.C3 is satisfied.
  • 5.
    An analogous statement holds true in the case of the up-and-out option control problem (see example 2.23, and [27]). Here $I$ is a singleton, $\mathbb{T}=\mathbb{R}_{+}$, and $\mathbb{A}^{i}=\mathbb{A}=\{0,1\}$ for the unique $i\in I$. Let $t\in\mathbb{R}_{+}$, $A_{<t}=\{1\}^{[0,t)}$,

### v2

@@ 4.3 Simple examples
For the examples from Subsection 2.3, both the basic version and its variation, we provide a reference choice structure, and a list of adapted choices, one for each exogenous information structure from Example 3.3.
@@
$\mathscr{C}=(\mathscr{C}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ defines a reference choice structure on ${\bm{\mathsf{X}}}$.
@@ 1. $c_{k\bullet}$ : $k\in\{1,2\}$ $c_{km},c_{\bullet m}$ : $k,m\in\{1,2\}$
2.(b) $c_{k\bullet}$ : $k\in\{1,2\}$ $c_{1g},c_{2m},c_{\bullet m}$ : $m\in\{1,2\},\,g\in M$
2.(c) $c_{k\bullet}$ : $k\in\{1,2\}$ $c_{1m},c_{2g},c_{\bullet m}$ : $m\in\{1,2\},\,g\in M$
@@ 3. $c_{f\bullet}$ : $f\in M$ $c_{kg},c_{\bullet g}$ : $k\in\{1,2\},\,g\in M$
$c^{\prime}_{kg}=$$\leavevmode\nobreak\ \{(\omega,k^{\prime},m^{\prime})\in W\mid k^{\prime}=k,\,m^{\prime}=g(\omega)\},$
@@
$\mathscr{C}^{\prime}=(\mathscr{C}^{\prime}_{{\bm{\mathsf{x}}}^{\prime}})_{{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime}}$ defines a reference choice structure on ${\bm{\mathsf{X}}}^{\prime}$.
@@ 3. $c^{\prime}_{f\bullet}$ : $f\in M$ $c^{\prime}_{kg},c^{\prime}_{\bullet g}$ : $k\in\{1,2\},\,g\in M$
Also note that a choice like $c_{kg}$ is conditional on the knowledge that ${\bm{\mathsf{x}}}_{k}$ has been chosen at the root of ${\bm{\mathsf{X}}}$, while $c_{\bullet g}$ is independent of the initial choice. This is reflected by the respective sets of immediate predecessors, see Lemma A.49. A similar remark is true for the variant, see Lemma A.50. Hence, as in [3, Section 5], choices can reflect the endogenous information, that is, the information about the position in the decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$, agents have. While $c_{kg}$ is a choice of perfect (endogenous) information, $c_{\bullet g}$ is not. The discussion on this theme will be continued in the second paper, see [40].
Concerning the sdf model $(F,\pi,{\bm{\mathsf{X}}})$ of the absent-minded driver story following Gilboa, with the exogenous information structure $\mathscr{F}$ as discussed in Subsection 3.3, let, for $i\in I=\{1,2\}$:
$\operatorname{Ex}_{i}=\underbrace{[\rho^{-1}(i)\times\{D\}]\cup[\rho^{-1}(3-i)\times\{H\}]\}}_{=\text{``exit''}},\qquad\operatorname{Ct}_{i}=\underbrace{[\rho^{-1}(i)\times\{H,M\}]\cup[\rho^{-1}(3-i)\times\{M\}]}_{\text{=``continue''}},$
and $\mathscr{C}_{{\bm{\mathsf{x}}}_{i}}^{i}=\{\operatorname{Ex}_{i},\operatorname{Ct}_{i}\}$. Futhermore, for any $E\in\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{i}}$, let
$c_{i}(E)=(W_{E}\cap\operatorname{Ex}_{i})\cup(W_{E^{\complement}}\cap\operatorname{Ct}_{i}),$
the choice of “agent” to exit in the event $E$ and to continue in the opposite event $E^{\complement}$. $E$ might be thought about as an event independent of $\rho$, allowing for individual “randomisation”. Let $C^{i}=\{c_{i}(E)\mid E\in\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{i}}\}$. That is, at both random moves ${\bm{\mathsf{x}}}_{i}$, $i\in I=\{1,2\}$, the active agent $i$ has two basic choices: “exit” and “continue”, between that $i$ can randomise depending on $i$’s exogenous information, that is, in an $\mathscr{F}_{{\bm{\mathsf{x}}}_{i}}^{i}$-measurable way. It is easily seen that $\mathscr{C}^{i}$ defines a reference choice structure on $\{{\bm{\mathsf{x}}}_{i}\}$ and that $C^{i}$ is a set of $\mathscr{F}^{i}$-$\mathscr{C}^{i}$-adapted choices, for both $i\in I$.
@@ 4.4 Action path stochastic decision forests
Finally, we consider the action path sdf $(F,\pi,{\bm{\mathsf{X}}})$ from Subsection 2.4, induced by action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$.
Let $t\in\mathbb{T}$. For any set $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and any family $A_{t}=(A_{t,\omega})_{\omega\in\Omega}\in\mathcal{P}(\mathbb{A})^{\Omega}$ of subsets of $\mathbb{A}$, let
$c(A_{<t},A_{t})=\{(\omega,f)\in W\mid f|_{[0,t)_{\mathbb{T}}}\in A_{<t},\,f(t)\in A_{t,\omega}\}.$
@@
So we consider choices that correspond to actions at a predefined time $t$. Again, there is some sort of duality here: action paths are the result of progressive choosing; choices are collections of action paths, essentially. More precisely, $c(A_{<t},A_{t})\in\mathscr{C}_{t}$ describes the choice of an action in $A_{t,\omega}$ in scenario $\omega$ and at time $t$, given the history is contained in $A_{<t}$. We assume that such an action is really possible for at least some scenario (Assumption AP.C0), that it really constitutes a choice in that there is an alternative (Assumption AP.C1), and that it is complete in the sense that it only trivially intersects with random moves (Assumption AP.C2).
@@
$\downarrow c=\{x_{u}(w)\mid w\in c,\,u\in\mathbb{T}\colon t<u\}\cup\{\{w\}\mid w\in c\}.$
@@
We fix an action index $i\in I$ (in perspective, modelling an agent), a set of random moves $\tilde{\bm{\mathsf{X}}}^{i}$ (of that agent), and an exogenous information structure $\mathscr{F}^{i}=(\mathscr{F}^{i}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}}$ (this agent is equipped with). We denote the canonical projection $\mathbb{A}\to\mathbb{A}^{i}$ by $p^{i}$. For all $t\in\mathbb{T}$ and ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, let $\mathscr{C}^{i}_{{\bm{\mathsf{x}}}}$ be a set of sets $c(A_{<t},A_{t})$ as above such that
@@
    $A_{t}=(A_{t,\omega})_{\omega\in\Omega}$ such that there is $A_{t}^{i}\in\mathscr{B}(\mathbb{A}^{i})$ satisfying, for all $\omega\in\Omega$,
@@
These properties are referenced as ($\mathscr{C}^{i}_{\bm{\mathsf{x}}}$.1) etc. Hence, a choice $c(A_{<t},A_{t})\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ allows for choosing a measurable set of individual actions for “agent” $i$ at the random move ${\bm{\mathsf{x}}}$ given the endogenous past $A_{<t}$.
@@
$\mathscr{C}^{i}=(\mathscr{C}^{i}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}}$ defines a reference choice structure on $\tilde{\bm{\mathsf{X}}}^{i}$.
Let $t\in\mathbb{T}$, $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$, $D\in\mathscr{E}$, and $g\colon D\to\mathbb{A}^{i}$. Let $A_{t}^{i,g}=(A_{t,\omega}^{i,g})_{\omega\in\Omega}$ be given by
@@
Let $c(A_{<t},i,g)=c(A_{<t},A_{t}^{i,g})$. Provided it is an element of $\mathscr{C}_{t}$, this models the choice of “agent” $i$, given an endogenous history in $A_{<t}$, to take the random action $g$, that is, the action $g(\omega)$ in scenario $\omega\in D$, at time $t$. The perfect (endogenous) information case corresponds to $A_{<t}=\{f|_{[0,t)_{\mathbb{T}}}\}$ for some $f\in\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f}\neq\emptyset$.
Example 4.43.
For $\mathbb{T}=\mathbb{R}_{+}$, singleton $\Omega$, singleton $A_{<t}=\{f|_{[0,t)_{\mathbb{T}}}\}$ for $t\in\mathbb{T}$ and $f\in\mathbb{A}^{\mathbb{T}}$, and $W=\Omega\times\mathbb{A}^{\mathbb{T}}\cong\mathbb{A}^{\mathbb{T}}$, $c(A_{<t},i,g)$ corresponds to the so-called “differential game” choice in [2, Section 2] (see also [5, Example 4.14]), where agent $i$ chooses the value of $g$ given the history $f|_{[0,t)_{\mathbb{T}}}$.
Example 4.44.
If $\mathbb{T}=\{0,1,\dots,N\}$ for some $N\in\mathbb{N}$ or $\mathbb{T}=\mathbb{N}$, then these $c(A_{<t},i,g)$ reflect the usual specification of choices in terms of adapted processes from decision problems in discrete time (see, e.g. [10]) if $g$ is supposed to be $\mathscr{F}_{\bm{\mathsf{x}}}^{i}$-adapted at all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $c(A_{<t},i,g)$ is available at. For the case of continuous-time, $\mathbb{T}=\mathbb{R}_{+}$ (see, e.g.  [38]), a similar remark can be made (note, however, that additional regularity conditions along time – progressive measurability, optionality, predictability, for example – are imposed for complete contingent plans of action, which is discussed in the third part).
@@
The next and final theorem of this paper is concerned with the following question, which can be motivated by the preceding Example 4.44. Provided $c=c(A_{<t},i,g)\in\mathscr{C}_{t}$, what is the link between the adaptedness of $c$ and the measurability of the function $g|_{D_{\bm{\mathsf{x}}}}$ with respect to $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$, for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $c$ is available at? [^18] As stated in the theorem below, the availability of $c$ at ${\bm{\mathsf{x}}}$ implies $D_{\bm{\mathsf{x}}}\subseteq D$. Indeed, in stochastic control and game theory, action path specifications are often implicitly used and choices are defined by assigning an action $g(\omega)$ to any scenario $\omega$, for $g$ measurable with respect to the value of the underlying filtration at the current time (see, e.g. [10, Chapter 8], [38, 14, 13]). Hence, if we can answer the question above by providing a tight link, then this customary modelling paradigm in stochastic control and game theory can be explained as a derivative of adapted choices in stochastic decision forests and can therefore be interpreted in terms of traditional decision theory.
@@
The theorem below affirms that the $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurability of $g|_{D_{\bm{\mathsf{x}}}}$ for relevant ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ is sufficient for the adaptedness of $c$. Moreover, it is also necessary, provided the reference choice structure sufficiently reflects the Borel $\sigma$-algebra on $\mathbb{A}^{i}$. For making this precise, let us call a set $\mathcal{M}$ stable under non-trivial intersections iff for all $A,B\in\mathcal{M}$ with $A\cap B\neq\emptyset$, we have $A\cap B\in\mathcal{M}$. This property is equivalent to saying that ${}\mathcal{M}\cup\{\emptyset\}$ is stable under intersections. For example, any partition has this property. Now, given some $t\in\mathbb{T}$, $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$, $i\in I$, $D\in\mathscr{E}$, and $g\colon D\to\mathbb{A}^{i}$ such that $\tilde{c}=c(A_{<t},i,g)\in\mathscr{C}_{t}$, we consider the following assumption:
@@
  • Assumption AP.C3 on the triple $(A_{<t},i,g)$ and on $\mathscr{C}^{i}$. For all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $\tilde{c}$ is available at, there is a generator $\mathscr{G}(\mathbb{A}^{i})$ of the Borel $\sigma$-algebra of $\mathbb{A}^{i}$, stable under non-trivial intersections, such that for all $G\in\mathscr{G}(\mathbb{A}^{i})$, upon letting $A^{i,G}_{t}=(A^{i,G}_{t,\omega})_{\omega\in\Omega}$ be given by $A^{i,G}_{t,\omega}=(p^{i})^{-1}(G)$ for $\omega\in D_{\bm{\mathsf{x}}}$ and $A^{i,G}_{t,\omega}=\emptyset$ for $\omega\notin D_{\bm{\mathsf{x}}}$, we have
    $c(A_{<t},A^{i,G}_{t})\in\mathscr{C}^{i}_{{\bm{\mathsf{x}}}}.$
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be the action path sdf induced by action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$. Further, let $i\in I$ be an action index, $\tilde{\bm{\mathsf{X}}}^{i}\subseteq{\bm{\mathsf{X}}}$ be a set of random moves, $\mathscr{F}^{i}$ be an exogenous information structure on $\tilde{\bm{\mathsf{X}}}^{i}$ and $\mathscr{C}^{i}$ be a reference choice structure on $\tilde{\bm{\mathsf{X}}}^{i}$ as above satisfying Axioms ($\mathscr{C}^{i}_{\bm{\mathsf{x}}}$.$k$), $k=1,\dots,4$. Let $t\in\mathbb{T}$, $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$, $D\in\mathscr{E}$, and $g\colon D\to\mathbb{A}^{i}$ such that $c(A_{<t},i,g)\in\mathscr{C}_{t}$.
Then, we have:
@@
    $c(A_{<t},i,g)$ is a non-redundant and $\tilde{\bm{\mathsf{X}}}^{i}$-complete choice.
@@
    For all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $c(A_{<t},i,g)$ is available at, we have $D_{\bm{\mathsf{x}}}\subseteq D$.
  3. 3.
    If for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $c(A_{<t},i,g)$ is available at $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}}$-measurable, then $c(A_{<t},i,g)$ is $\mathscr{F}^{i}$-$\mathscr{C}^{i}$-adapted.
  4. 4.
    If Assumption AP.C3 is satisfied for $(A_{<t},i,g)$ and $\mathscr{C}^{i}$, and if moreover $c(A_{<t},i,g)$ is $\mathscr{F}^{i}$-$\mathscr{C}^{i}$-adapted, then $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}}$-measurable for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $c(A_{<t},i,g)$ is available at.
@@
Recall that we consider the action path sdf $(F,\pi,{\bm{\mathsf{X}}})$ induced by action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$.
@@
  • Suppose that $W=\Omega\times\mathbb{A}^{\mathbb{T}}$. Fix some $i\in I$ and suppose that $\mathbb{A}^{i}$ has at least two elements. Let $t\in\mathbb{T}$ and $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$. Then, for any map $g\colon\Omega\to\mathbb{A}^{i}$ we have $c(A_{<t},i,g)\in\mathscr{C}_{t}$, that is, it satisfies Assumptions AP.C0, AP.C1, and AP.C2. Moreover, if for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ contains all $c(A_{<t},A_{t})$ ranging over all $A_{t}$ satisfying ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.$k$), $k=2,3,4$, then Assumption AP.C3 is satisfied for $(A_{<t},i,g)$ and $\mathscr{C}^{i}$.
  • In case of the timing problem (see Example 2.26, and, e.g. [47, 33]), we have $\mathbb{A}^{i}=\{0,1\}$ for all $i\in I$. Fix $i\in I$. Let $t\in\mathbb{T}$ and $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ be a non-empty set of componentwise decreasing paths $f_{t}$ such that $p^{i}\circ f_{t}=1_{[0,t)_{\mathbb{T}}}$. Then, for all $g\colon\Omega\to\{0,1\}$, we have $c(A_{<t},i,g)\in\mathscr{C}_{t}$, that is, it satisfies Assumptions AP.C0, AP.C1, AP.C2. Moreover, if for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ contains all $c(A_{<t},A_{t})$ ranging over all $A_{t}$ satisfying ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.$k$), $k=2,3,4$, then Assumption AP.C3 is satisfied for $(A_{<t},i,g)$ and $\mathscr{C}^{i}$.
  • An analogous statement holds true in the case of the up-and-out option control problem (see Example 2.26, and [31]). Here $I$ is a singleton, $\mathbb{T}=\mathbb{R}_{+}$, and $\mathbb{A}^{i}=\mathbb{A}=\{0,1\}$ for the unique $i\in I$. Let $t\in\mathbb{R}_{+}$ such that, with